OpenStudy (safia21):

Solve proportion Latanya made a 17 minute phone call from her hotel in France and was charged 17 euro. At the time, $1 worth 0.82 euro. Find the cost per minute of the calls in dollars

OpenStudy (anonymous):

This is a conversion type of question

OpenStudy (anonymous):

So.. it tells you $1 = 0.82 euro 1/ 0.82= ??? <---This will be your conversion factor Since it costed you 17 euros, how many dollars would it be? 17 x ??? = $ ___

OpenStudy (safia21):

i dont know

OpenStudy (anonymous):

Could you at least try to plug in the numbers into a calculator? I already set up the equations.

OpenStudy (safia21):

im consfused so its 1/0.82 = 17 times ???

OpenStudy (anonymous):

Okay forst of all do you know how to convert things?

OpenStudy (anonymous):

first*

OpenStudy (safia21):

yes

OpenStudy (anonymous):

Okay so if 1 full dollar = 0.82 of a dollar, can we agree that an euro is worth more than a dollar?

OpenStudy (safia21):

yes

OpenStudy (anonymous):

Alright so if we want to know how many dollars 17 euro equals? It would help if we knew how much 1 full euro equals to, yes?

OpenStudy (safia21):

yes

OpenStudy (anonymous):

But let's say we want to know by how many times more a euro is worth, than a dollar. So 1/0.82 is...?

OpenStudy (safia21):

you want me to divide that

OpenStudy (anonymous):

Mhmm

OpenStudy (safia21):

what?

OpenStudy (anonymous):

Yes divide it.

OpenStudy (safia21):

1.213

OpenStudy (anonymous):

Alright so let's just round it to 2 decimal places. 1.21 <-- this is how much one euro is worth in dollars

OpenStudy (anonymous):

So now that you know what one euro is worth, how much would 17 euros be worth?

OpenStudy (anonymous):

Did you get it?

OpenStudy (gw2011):

1/0.82=x/17 Cross multiply to get: 0.82x=17 x=$20.73 Divide: $20.73/17 min=$1.22 per minute